OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = x +

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ignore me

OpenStudy (igreen):

First we write the equation in point-slope form, then we simplify it. \(y - y_1 = m(x - x_1)\) Where (x1, y1) is the given point, and 'm' is the slope. Can you plug in (6, 2) with a slope of 3 into that equation? @gummybear1021

OpenStudy (igreen):

Vertex form is for parabolas..lol. @swagmaster47

OpenStudy (anonymous):

Oopsie

OpenStudy (anonymous):

umm y=4x+2?

OpenStudy (anonymous):

idk Im confused

OpenStudy (igreen):

Well first we plug in the points into the equation: \(x_1 = 6\) \(y_1 = 2\) \(m = 3\) \(y - y_1 = m(x - x_1)\) \(y - \color{red}2 = \color{red}3(x - \color{red}6)\) Okay, now that we have our equation in point-slope form, all we have to do is simplify to get it into slope-intercept form. So can you simplify this part? \(3(x - 6)\)

OpenStudy (anonymous):

umm idk, ummmm 3?

OpenStudy (igreen):

It's not going to be a 1 number answer.. \(3(x - 6) \rightarrow (3 \times x) + (3 \times -6)\) So can you multiply 3 * x and 3 * -6?

OpenStudy (anonymous):

but whats x?

OpenStudy (anonymous):

You dont need to find x, it is a function :)

OpenStudy (anonymous):

Unless you want to find the x intercepts

OpenStudy (anonymous):

im so confused...

OpenStudy (igreen):

Okay, can you tell me what 3 * x is?

OpenStudy (anonymous):

3x right?

OpenStudy (igreen):

Yes! Now can you tell me what 3 * -6 is?

OpenStudy (anonymous):

-18

OpenStudy (igreen):

Yes, so \(3(x - 6)\) simplifies to \(3x - 18\). So now we have: \(y - 2 = 3x - 18\) Our last step to converting into slope-intercept form is the add 2 to both sides. And what's -18 + 2?

OpenStudy (anonymous):

-16?

OpenStudy (igreen):

Yep, so our final answer is: \(y = 3x - 16\)

OpenStudy (anonymous):

thnx